A car travels 64 km in 2 hours.What distance will it travel in 6(1)/(4) hours.If the speed remains the same.
Answer by oscargut(2103)

64 km in 2 hs
speed = 64/2 = 32 km/h
it travels 6(1/4) in [6+1/4]32 hs = 32(6)+8 = 200
Answer: 200km
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com